A stock has an expected return of 11.2 percent, its beta is 0.4, and the risk-free rate is 3.7 percent. What must the expected return on the market be? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)
Answer #1

Calculation of Expected return on market ( Market return )

As per CAPM,

Expected return of stock = Risk free rate + Beta * (Market return - Risk free rate)

11.2% = 3.7% + .4(Market return -3.7)

11.2-3.7 = .4(Market return -3.7)

7.5 = .4(Market return -3.7)

7.5/.4 = Market return -3.7

18.75 = Market return -3.7

So Market return =18.75 + 3.7 =22.45%